OpenStudy (raheelafzaal):

solve the integration f(sinx+cosx)^2dx

OpenStudy (anonymous):

is this \[\int(\sin(x)+\cos(x))^2dx\]?

OpenStudy (raheelafzaal):

yes

OpenStudy (anonymous):

square and get \[\sin^2(x)+\cos^2(x)+2\sin(x)\cos(x)=1+\sin(2x)\] then it should be ok

OpenStudy (anonymous):

\[\int 1dx+\int\sin(2x)dx\]

OpenStudy (raheelafzaal):

can u tell me easy method to sove like this integrations?

OpenStudy (anonymous):

not sure what you mean. the original integral wants you to square.

OpenStudy (raheelafzaal):

ok

OpenStudy (anonymous):

if you mean how to solve \[\int\sin(2x)dx\] you visualize what you would get if you take the derivative of \[\cos(2x)\] you would get \[-2\sin(2x)\] but you want just \[\sin(2x)\] so your integral is \[-\frac{1}{2}\cos(2x)\]

OpenStudy (raheelafzaal):

ok ok

OpenStudy (raheelafzaal):

thanks my dear

OpenStudy (raheelafzaal):

\[\int\limits_{\sec^2xcosec^2x}^{dx}\]

OpenStudy (raheelafzaal):